OpenStudy (anonymous):

Long division (12x^2-5x-2) divided by (3x-2) i know the answer to the problem but i need a step by step explanation on how to get the answer!! Please help thank you!

OpenStudy (anonymous):



OpenStudy (anonymous):

how did you 3x? wouldnt it be -13x?

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

-5x-(-8x)=-5x+8x=3x or simply change the sign and add.

OpenStudy (anonymous):

oh okay thank you!!! i thought it was -5+(-8) = -13 @surjithayer

OpenStudy (anonymous):

yw